I specifically need help with part c and d Problem #1 (10 points) Customers arrive to a fast food restaurant with two servers with interarrival times that follow an exp
Answer #1

The M/M/2 Queue The underlying CTMC for the number of customers in the queue has a stationary distribution if and only if p<1

b)

1+1 01 1-

Pa

c is number of servers

= (1/5 )/(2 * 1/8)

= 0.8

hence

pi_0 = (1 - 0.8)/(1 + 0.8) = 0.1111

c)

server utilization = \rho = 0.8

d)

Lq = 2 * 0.8^3 / (1- 0.8^2) =

2.8444

e)

Wq = 0.237

Please give me a thumbs-up if this helps you out. Thank you! :)